# Multimodal Task Vectors Enable Many-Shot Multimodal In-Context Learning

### v1

@@ Abstract
The recent success of interleaved Large Multimodal Models (LMMs) in few-shot learning suggests that in-context learning (ICL) with many examples can be promising for learning new tasks. However, this many-shot multimodal ICL setting has one crucial problem: it is fundamentally limited by the model’s context length set at pretraining. The problem is especially prominent in the multimodal domain, which processes both text and images, requiring additional tokens. This motivates the need for a multimodal method to compress many shots into fewer tokens without finetuning. In this work, we enable LMMs to perform multimodal, many-shot in-context learning by leveraging Multimodal Task Vectors (MTV)—compact implicit representations of in-context examples compressed in the model’s attention heads. Specifically, we first demonstrate the existence of such MTV in LMMs and then leverage these extracted MTV to enable many-shot in-context learning for various vision-and-language tasks. Our experiments suggest that MTV can scale in performance with the number of compressed shots and generalize to similar out-of-domain tasks without additional context length for inference.
footnotetext: *Denotes Equal Contribution
@@ 1 Introduction
Figure 1: Overview of Multimodal Task Vectors (MTV). In the standard multimodal in-context learning (ICL) paradigm (top), the number of shots is limited by an LMM’s context length. We solve this issue by first calculating the mean activations of the last token’s attention heads (grey squares in the LMM) given a set of multimodal ICL examples and then finding a set of attention head locations (yellow squares in the LMM) that best align with the downstream task. These mean activations are then replaced directly in these attention heads locations, enabling many-shot multimodal ICL.
@@
Our method can be described in three steps. First, given a set of many-shot multimodal ICL examples, we calculate the mean activations of the attention heads across multiple inference iterations. Second, to avoid the context length constraint, we select a set of attention heads in the model to store the mean activations of the ICL examples. However, since the downstream task may be zero-shot or use a different number of ICL examples, we select a set of examples aligned with its form. We then use these examples to find an optimal set of LMM head locations where the many-shot examples will be encoded. We refer to these mean activations and locations as MTV, which implicitly encodes the many-shot multimodal examples for use in the downstream task. Finally, for downstream inference, we replace the mean activations from Step 1 with the attention head locations found in Step 2. Since we input examples to the LMM across different iterations in Step 1, Multimodal Task Vectors can implicitly encode more examples than are allowable by the context limit. We find that utilizing many examples for extracting MTV surpasses performance on zero-shot and most standard few-shot ICL settings, suggesting the effectiveness of our method. Another key benefit of our method is that it frees up tokens for the model during downstream inference compared to standard few-shot ICL methods.
@@ 2 Related Works
However, scaling ICL in text-only LLMs is a challenge due to the intrinsic context length. One method to increase context length in these models is to apply positional interpolation methods [10, 62]. However, research on these longer-context models finds that they struggle to use the entire context for ICL [51, 45]. Moreover, as inference on long contexts of inputs is also time and memory-expensive, it is unclear whether simply scaling the context of models is practical for enabling multimodal many-shot ICL. This has led to work that looks to compress explicit input tokens [20, 75, 11, 57, 71, 35]. But crucially, many of these methods require finetuning and only try to preserve performance. Our work is different in that it is the first to enable multimodal models with many-shot ICL capabilities, while also having the benefit of improving on complex VL tasks without finetuning.
@@ 3 Multimodal Task Vectors
To address the challenge of performing many-shot multimodal in-context learning, we demonstrate the existence of MTV in LMMs and then leverage them for many-shot multimodal ICL. We begin by describing some background on multimodal ICL and task vectors (Section 3.1). We then introduce our three-step approach: (i) We calculate the mean activations of the attention heads from the many-shot multimodal ICL examples (Section 3.2); (ii) We then extract the set of LMM attention heads locations that best align to the downstream task using an adapted version of the REINFORCE [89] algorithm (Section 3.3); and (iii) We replace the calculated mean activation values into the LMM for a downstream task (Section 3.4). The method overview is also illustrated Figure 1.
@@ 3.1 Preliminaries
We start with a background on task vectors for some task $j$. Given a model $F$, we denote the set of attention-head locations as $\lambda=\{l\mid\forall l\in F\}$ where each location $l$ is indexed as $l=(h,m)$ for the $h^{\mathrm{th}}$ layer and $m^{\mathrm{th}}$ attention head. We define the task vectors as follows: (1) the task vector values $\mu_{j}$ are a subset of mean activations produced by the attention heads of $F$ given examples of a task, and (2) the task vector locations $\lambda_{j}$, which denotes a subset of the attention head indices per task. Thus, the task vector is $(\mu_{j},\lambda_{j})$. For inference, $\mu_{j}$ replaces the activation values of the heads in $\lambda_{j}$.
@@ 3.3 Step 2: Extract MTV Attention Head Locations
From these examples, we utilize an adapted version of the REINFORCE [89] algorithm—an iterative policy optimization method that can be used to find task vector locations. Given an LMM $F$, we first select a proposed set of attention head locations by sampling a Bernoulli distribution over the locations multiple times. Next, we directly replace the values of the selected attention heads with the corresponding mean activations $\mu_{l,j}$. Then, after prompting the model with the query $Q_{s}$, we use the negative cross-entropy loss between the LMM’s output logits and the logits of the ground-truth response $R_{s}$ to optimize the Bernoulli distribution. By optimizing the Bernoulli distribution across $S$ iterations, we are finding the best attention head locations $\lambda^{\mathrm{MTV}}_{j}$ for patching in our mean activations. Finally, we can extract $\lambda^{\mathrm{MTV}}_{j}$, the optimized indices of attention heads, by sampling our optimized Bernoulli distribution.
@@
It is important to note that MTV_EXTRACT does not require finetuning of the LMM parameters, but rather only inference calls. We describe further the underlying details of our adapted MTV_EXTRACT algorithm in Section B of the Supplementary. Having found $\lambda^{\mathrm{MTV}}_{j}$ and $\mu_{l,j}$, we describe in what follows, the final procedure to use MTV for inference.
@@ 4.1 Implementation Details
We implemented MTV using PyTorch [60]. We used each model’s respective official implementation. While the compute and memory requirements differ slightly between models, all our experiments can be run on a single NVIDIA A6000 GPU. For additional information, refer to Supplementary Section C. Our model and weights will be released upon acceptance, and our code is in Supplementary.
@@ 4.2 Models
In this work, we apply MTV to the following interleaved LMMs as they are better-suited for multimodal ICL as shown by [16]: (1) QwenVL [6] is a LLaMA-based model that has the ability to process high-resolution images, and its two-stage pre-training methodology, which includes multi-task finetuning and interleaved text-image data. (2) Idefics2-8B [39] is a Mistral-based model that benefits from its pre-training on the expansive OBELICS dataset, which comprises a web-scale collection of interleaved image-text documents. We utilize the base version of the model. This demonstrates multimodal in-context learning abilities. (3) ViLA-1.5-8B. ViLA-1.5-8B [47] is an architecture that leverages LLaMA-3 as the LLM backbone. As in others, a significant portion of the model’s pretraining data is interleaved text-image data.
@@ 4.3 Datasets
We briefly describe the tasks and datasets we evaluate our method on. More details about the datasets and their setup can be found in Section C.
@@ 5.3 MTV heads generalize to other tasks
5.4 LoRA Finetuning as an upper bound
In Table 2(b), we compare our method to LoRA finetuning [29]. To do this, we use LoRA on the same number of examples as MTV uses from the train set and evaluate not only on the validation set but also on the validation set of another similar dataset. In particular, for a ViLA-1.5-8B model finetuned on VizWiz, we report accuracy on both VizWiz and OK-VQA validation sets. It can be seen that LoRA finetuning is indeed an upper bound on the dataset the model was finetuned on. However, we show that LoRA leads to overfitting on the finetuned dataset and forgetting the zero-shot capabilities. In contrast, we also show that MTV not only improves zero-shot capabilities but can generalize to similar tasks with only a few inference examples Table 1(b) and Table 2(a).
@@ 5.5 Comparison to other methods
We compare our method to two different methods that can find task vectors: Visual Task Vectors (VTV) [27] and Function Vectors (FV) [79]. Originally, these works could not be applied as-is to support multimodal ICL, but here, we have implemented a version that follows the original exactly with only minor modifications to allow performing our evaluated multimodal tasks. More details about the methods can be found in Section B in the Supplementary. In our experiments Table 2(b), we find that MTV surpasses both methods on VizWiz and OK-VQA. VTV are image-only task vectors that use only one-shot image examples for fixed small $T$ iterations, and they calculate the mean activations and the locations together without aligning to the downstream task. FV are text-only task vectors that use Causal Mediation Analysis [61] to extract task vector locations. But then a vector of mean activations at these locations is simply added after the output of an arbitrary layer. The results suggest the importance of finding the task vectors by decoupling the calculation of the mean activations and locations in two separate steps to perform many-shot multimodal ICL more effectively for complex multimodal tasks.
@@ 7 Limitations
Acknowledgements.
We would like to thank Suzie Petryk and Grace Luo for helpful feedback and discussions. This project was supported in part by DoD, including PTG and/or LwLL programs, as well as BAIR’s industrial alliance programs.
@@ A.1 Additional Experiments
Here we provide additional ablations that further illustrate different characteristics of MTV.
@@
MTV one-to-one comparison with ICL Table 4(b). Although not directly comparable, we consider an extreme case of MTV where we encode only 4-shots of ICL examples for 1 iteration. This matches the exact setting used in standard 4-shot ICL. Interestingly, MTV applied to both VizWiz and OK-VQA exceeds performance on the 4-shot-ICL case and even MTV formulated on 4-shots per 100 iterations for calculating the mean activations. This result suggests that there may be scope for MTV to be effective in both high and low-data regimens. More research needs to be done to explore this idea.
Effect of permutation order of examples. We consider applying five random seeds to both 4-shot-ICL and MTV extracted on 4-shots per 100 iterations on VizWiz. We find the 4-shot-ICL average and standard deviation to be 41.3 % ($\pm$ .8%) and the MTV average and standard deviation to be 45.2 % ($\pm$ .7 %). This suggests that MTV is stable across different permutations of the given ICL examples.
Scaling on Flowers Dataset.
@@
We provide additional results on the scaling property of MTV on the Flowers dataset. We again note that the examples are 2-way, one-shot examples with 2 examples (one positive and one negative) for each sample. As in the main paper, we fix 1 shot per iteration to calculate the mean activations, scaling up to 500 total examples used. Our results show that there is a saturation of MTV at 100 examples (i.e., 1 example per 100 iterations). While this still indicates some scaling as the result is an improvement over 20 examples, the results show that the task vector can reach its best accuracy with fewer shots depending on the complexity of the task. Future work to probe more deeply into the scaling nature of MTV across different tasks would be valuable.
Appendix B Additional Method Details

### v2

@@ Abstract
The recent success of interleaved Large Multimodal Models (LMMs) in few-shot learning suggests that in-context learning (ICL) with many examples can be promising for learning new tasks. However, this many-shot multimodal ICL setting has one crucial problem: it is fundamentally limited by the model’s context length set at pretraining. The problem is especially prominent in the multimodal domain, which processes both text and images, requiring additional tokens. This motivates the need for a multimodal method to compress many shots into fewer tokens without finetuning. In this work, we enable LMMs to perform multimodal, many-shot in-context learning by leveraging Multimodal Task Vectors (MTV)—compact implicit representations of in-context examples compressed in the model’s attention heads. Specifically, we first demonstrate the existence of such MTV in LMMs and then leverage these extracted MTV to enable many-shot in-context learning for various vision-and-language tasks. Our experiments suggest that MTV can scale in performance with the number of compressed shots and generalize to similar out-of-domain tasks without additional context length for inference. Code: https://github.com/Brandon3964/MultiModal-Task-Vector
@@ 1 Introduction
Our method can be described in three steps. First, given a set of many-shot multimodal ICL examples, we calculate the mean activations corresponding to the last token across multiple inference iterations. Second, to avoid the context length constraint, we select a set of attention heads in the model to store the mean activations of the ICL examples. However, since the downstream task may be zero-shot or use a different number of ICL examples, we select a set of examples aligned with its form. We then use these examples to find an optimal set of LMM head locations where the many-shot examples will be encoded. We refer to these mean activations and locations as MTV, which implicitly encodes the many-shot multimodal examples for use in the downstream task. Finally, for downstream inference, we replace the mean activations from Step 1 with the attention head locations found in Step 2. Since we input examples to the LMM across different iterations in Step 1, Multimodal Task Vectors can implicitly encode more examples than are allowable by the context limit. We find that utilizing many examples for extracting MTV surpasses performance on zero-shot and most standard few-shot ICL settings, suggesting the effectiveness of our method. Another key benefit of our method is that it frees up tokens for the model during downstream inference compared to standard few-shot ICL methods. An overview of our method is shown in Figure 1.
@@
Figure 1: Multimodal Task Vectors (MTV) Overview. We overcome an LMM’s context length limitation by encoding many shots of multimodal examples as activations in the LMM’s latent space. We then directly replace this encoding into the LMM’s activation space during downstream inference.
@@ 2 Related Works
However, scaling ICL in text-only LLMs is a challenge due to the intrinsic context length. One method to increase context length in these models is to apply positional interpolation methods [10, 63]. However, research on these longer-context models finds that they struggle to use the entire context for ICL [51, 45]. Moreover, as inference on long contexts of inputs is also time and memory-expensive, it is unclear whether simply scaling the context of models is practical for enabling multimodal many-shot ICL in open-source models. There is some early evidence of multimodal many-shot ICL being effective in closed-source models [35], so the question arises as to how to achieve something similar for open-source models. This has led to work that looks to compress explicit input tokens [20, 77, 11, 58, 73, 34]. But crucially, many of these methods require finetuning and only try to preserve performance. Our work is different in that it is the first to enable multimodal models with many-shot ICL capabilities, while also improving on complex VL tasks without finetuning.
@@ 3 Multimodal Task Vectors
Figure 2: Multimodal Task Vectors (MTV). In the standard multimodal in-context learning (ICL) paradigm, the number of shots is limited by an LMM’s context length. We solve this issue by first finding the mean activations corresponding to the last token of the examples’ input (Step 1), and then calculating a set of attention head locations (Step 2) that best align with the downstream task. These mean activations are then replaced directly in these attention head locations (Step 3), enabling many-shot multimodal ICL.
To address the challenge of performing many-shot multimodal in-context learning, we demonstrate the existence of MTV in LMMs and then leverage them for many-shot multimodal ICL. We begin by describing some background on multimodal ICL and task vectors (Section 3.1). We then introduce our three-step approach: (i) We calculate the mean activations of the attention heads from the many-shot multimodal ICL examples (Section 3.2); (ii) We then extract the set of LMM attention heads locations that best align to the downstream task using an adapted version of the REINFORCE [91] algorithm (Section 3.3); and (iii) We replace the calculated mean activation values into the LMM for a downstream task (Section 3.4). The detailed method visual is shown in Figure 1.
@@ 3.1 Preliminaries
We start with a background on task vectors for some task $j$. Given a model $F$, we denote the set of attention-head locations as $\lambda=\{l\mid\forall l\in F\}$ where each location $l$ is indexed as $l=(h,m)$ for the $h^{\mathrm{th}}$ layer and $m^{\mathrm{th}}$ attention head. Now, task vectors utilize the intermediate outputs of an LMM, called activations. For a given input sequence of written in terms of its tokens $x=\{x_{1},x_{2},\ldots,x_{T}\}$, each attention head $(h,m)$ produces an activation $z_{l}\in\mathbb{R}^{\frac{d}{H}}$ for each token $x_{i}$, where $d$ is the model’s embedding dimension and $H$ is the number of heads. These activations are simply the output vectors of each attention head before any linear projection. While each head’s activation is typically concatenated with others and projected to form the layer’s output, task vectors specifically utilize the pre-projection activations of the final token $x_{T}$ from each attention head. We thus define the task vectors as follows: (1) the task vector values $\mu_{j}$ are a subset of mean activations produced by the attention heads of $F$ given examples of a task, and (2) the task vector locations $\lambda_{j}$, which denotes a subset of the attention head indices per task. Thus, the task vector is $(\mu_{j},\lambda_{j})$. For inference, $\mu_{j}$ replaces the activation values of the heads in the locations given by $\lambda_{j}$.
@@ 3.3 Step 2: Extract MTV Attention Head Locations
From these examples, we utilize an adapted version of the REINFORCE [91] algorithm—an iterative policy optimization method that can be used to find task vector locations [27]. Given an LMM $F$, we first select a proposed set of attention head locations by sampling a Bernoulli distribution over the locations multiple times. Next, we directly replace the values of the selected attention heads with the corresponding mean activations $\mu_{l,j}$. Then, after prompting the model with the query $Q_{s}$, we use the negative cross-entropy loss between the LMM’s output logits and the logits of the ground-truth response $R_{s}$ to optimize the Bernoulli distribution. By optimizing the Bernoulli distribution across $S$ iterations, we are finding the best attention head locations $\lambda^{\mathrm{MTV}}_{j}$ for patching in our mean activations. Finally, we can extract $\lambda^{\mathrm{MTV}}_{j}$, the optimized indices of attention heads, by sampling our optimized Bernoulli distribution.
@@
It is important to note that MTV_EXTRACT does not require finetuning of the LMM parameters, but rather only inference calls. We describe further the underlying details of our adapted MTV_EXTRACT algorithm in Section A.2 of the Supplementary. Having found $\lambda^{\mathrm{MTV}}_{j}$ and $\mu_{l,j}$, we describe in what follows, the final procedure to use MTV for inference.
@@ 4.1 Implementation Details
We implemented MTV using PyTorch [61]. We used each model’s respective official implementation. While the compute and memory requirements differ slightly between models, all our experiments can be run on a single NVIDIA A6000 GPU. For additional information, refer to Supplementary Section B. Our model and weights will be released upon acceptance, and our code is in Supplementary.
@@ 4.2 Models
In this work, we apply MTV to the following interleaved LMMs as they are better-suited for multimodal ICL as shown by [16]: (1) QwenVL [6] is a LLaMA-based model that has the ability to process high-resolution images, and its two-stage pre-training methodology, which includes multi-task finetuning and interleaved text-image data. (2) Idefics2-8B [39] is a Mistral-based model that benefits from its pre-training on the expansive OBELICS dataset, which comprises a web-scale collection of interleaved image-text documents. We utilize the base version of the model. This demonstrates multimodal in-context learning abilities. (3) LLaMA3-ViLA-1.5-8B (abbreviated as VILA-1.5-8B). ViLA-1.5-8B [47] is an architecture that leverages LLaMA-3 as the LLM backbone. As in others, a significant portion of the model’s pretraining data is interleaved text-image data. (4) MANTIS-LLaMA3-8B. MANTIS-LLaMA3-8B [33] is a combination of a SigLIP [98] visual encoder and LLaMA3 [55] language model finetuned using the MANTIS dataset, a specially curated multi-image dataset that emphasizes co-reference, reasoning, comparing, temporal understanding.
We show the number of tokens per image embedding for each model in Table 1 to illustrate the especial importance of MTVs in the image-text domain:
Table 1: Per Image Embedding Token Length and Total Context Length for Models
Model Name Per Image Token Length Total Context Length
VILA-1.5-8B 144 8192
Idefics2-8B 64 8192
QwenVL 256 8192
MANTIS-LLaMA3-8B 64 8192
@@ 4.3 Datasets
We briefly describe the tasks and datasets we evaluate our method on. More details about the datasets and their setup can be found in Section B.
@@ 5.3 MTV heads generalize to other tasks
5.4 Finetuning as an upper bound
In Table 3(b), we compare our method to finetuning. To do this, we use finetune on the same number of examples as MTV uses from the train set and evaluate not only on the validation set but also on the validation set of another similar dataset. In particular, for a ViLA-1.5-8B model finetuned on VizWiz, we report accuracy on both VizWiz and OK-VQA validation sets. It can be seen that finetuning is indeed an upper bound on the dataset the model was finetuned on. However, we show that finetuning leads to overfitting on the finetuned dataset and even forgetting the zero-shot capabilities. In contrast, we also show that MTV not only improves zero-shot capabilities but can generalize to similar tasks with only a few inference examples Table 2(b) and Table 3(a).
@@ 5.5 Comparison to other methods
We compare our method to two different methods that can find task vectors: Visual Task Vectors (VTV) [27] and Function Vectors (FV) [81]. Originally, these works could not be applied as-is to support multimodal ICL, but here, we have implemented a version that follows the original exactly with only minor modifications to allow performing our evaluated multimodal tasks. More details about the methods can be found in Section A.2 in the Supplementary. In our experiments Table 3(b), we find that MTV surpasses both methods on VizWiz and OK-VQA. VTV are image-only task vectors that use only one-shot image examples for fixed small $T$ iterations, and they calculate the mean activations and the locations together without aligning to the downstream task. FV are text-only task vectors that use Causal Mediation Analysis [62] to extract task vector locations from only the output activations of the last token. The results suggest the importance of finding the task vectors by decoupling the calculation of the mean activations and locations in two separate steps to perform many-shot multimodal ICL more effectively for complex multimodal tasks.
@@ 7 Limitations
8 Acknowledgements
We would like to thank Deva Ramanan, Grace Luo, and Suzanne Petryk for their insightful feedback and discussions. This project has received funding from Prof. Darrell’s group in part by DoD, including PTG and/or LwLL programs, as well as BAIR’s industrial alliance programs.
@@ A.1 Additional Experiments
Here we provide additional experiments and ablations that further illustrate different characteristics of MTV.
Motivation for encoding shots in the activation space.. We highlight our paper’s motivation in addressing the context length token limitation of LMMs by encoding ICL shots in the activation space. An additional limiting factor in the token space is the physical constraints of memory and runtime, which we ablated in Section 5.6 of the paper. For example, 25-shot ICL is actually the maximum number of vanilla ICL shots that can be run on a single 48GB A6000 GPU for Qwen-VL. We demonstrate the degradation of increasing numbers of multimodal token-space ICL shots (VizWiz-QwenVL) in Table 5(a).
Effect of shot quality on MTV:. We assess the connection between textual and activation-space shot quality by comparing MTV using random selection with MTV using high-quality shots selected with the Facility Location algorithm [70]. We apply MTV to QwenVL and use the Qwen GTE embedding model to obtain embeddings for the Facility Location algorithm and present the results in Table 5(b). Excitingly, we find that high-quality shots do indeed lead to significant improvements in MTV performance.
MTV with noisy exemplars.. We compare the robustness of MTV compared to that of vanilla ICL. For QwenVL on VizWiz and OKVQA, we replace 1 of the 4 examples in each iteration of 4-shot ICL and 4-shot-100-iteration MTV with an example from the opposite dataset. We report both accuracy and degradation in Table 5(c)
Table 5: ICL Degradation, Shot-Quality Impact, and Stability (Left) Degradation of ICL with increasing number of shots. (Right) Impact of shot quality on MTV and stability of ICL vs MTV with noisy examples.
(a) ICL Degradation with Increasing Shots
ICL Shots Acc. % Acc. Increase
0 35.2 -
4 42.0 6.8
8 44.3 2.3
16 46.9 2.6
20 49.0 2.1
25 49.8 0.8
(b) MTV with High-Quality Shots
Model VizWiz
QwenVL-7B 35.2
+ MTV 45.6
+ MTV + F.L. Shots 58.1
(c) Stability of ICL vs MTV using QwenVL
VizWiz OK-VQA
4-shot ICL 41.0 (-1.0) 61.5 (-0.5)
MTV 43.4 (-2.2) 61.9 (-0.1)
@@
MTV one-to-one comparison with ICL -  Table 6(b). Although not directly comparable, we consider an extreme case of MTV where we encode only 4-shots of ICL examples for 1 iteration. This matches the exact setting used in standard 4-shot ICL. Interestingly, MTV applied to both VizWiz and OK-VQA exceeds performance on the 4-shot-ICL case and even MTV formulated on 4-shots per 100 iterations for calculating the mean activations. This result suggests that there may be scope for MTV to be effective in both high and low-data regimens. More research needs to be done to explore this idea.
@@
Scaling on Flowers Dataset. We provide additional results on the scaling property of MTV on the Flowers dataset. We again note that the examples are 2-way, one-shot examples with 2 examples (one positive and one negative) for each sample. As in the main paper, we fix 1 shot per iteration to calculate the mean activations, scaling up to 500 total examples used. Our results show that there is a saturation of MTV at 100 examples (i.e., 1 example per 100 iterations). While this still indicates some scaling as the result is an improvement over 20 examples, the results show that the task vector can reach its best accuracy with fewer shots depending on the complexity of the task. Future work to probe more deeply into the scaling nature of MTV across different tasks would be valuable.
MTV at extreme shot counts. We delve further into the scaling capabilities by evaluating the performance of MTV at the maximum number of VizWiz shots per iteration allowable by the memory constraints of a single NVIDIA RTX A6000. The experiment shown in Table 7(a) indicates that while MTV does continue to scale, there is also certainly a saturation point for VizWiz. The exact saturation point likely depends on the specific task.
Effect of permutation. We consider applying five random seeds to different configurations of MTV comparing its variability under permutation of example order to standard few-shot ICL. We present the mean and standard deviation for these experiments in  Table 7(b). Although not significant, in both 4-shot and 8-shot settings, MTV shows less variability to example permutation. This intuitively makes sense as more examples are averaged over multiple iterations, leading to more stable performance across different seeds.
Table 7: QwenVL-7B: Generalization & Variability to Permutation (Left) Evaluation of MTV on extreme shot-iteration counts. (Right) Variability to permutation across 5 seeds.
(a) QwenVL-7B: Extreme shot-count performance
Model Accuracy (%)
QwenVL-7B
    + MTV_20shot_100it 54.9
    + MTV_20shot_200it 55.1
    + MTV_25shot_100it 56.4
    + MTV_25shot_200it 51.4
(b) Permutation Variability
Model Accuracy (%)
QwenVL-7B
    + MTV_4-shot_100it 45.2 (0.7)
    + MTV_4-shot_200it 48.3 (0.4)
    + MTV_8-shot_100it 50.4 (0.9)
    + MTV_8-shot_200it 51.8 (0.6)
    + 4-shot ICL 41.3 (0.8)
    + 8-shot ICL 42.9 (1.5)
MTV for language-only tasks. While we show the importance of MTV especially for vision-language tasks, the methodology can be a powerful way to learn tasks in the language-only domain as well. We demonstrate in Table 8(a) the effectiveness of MTV on two common LLM tasks using LLaMA-3-8B [55]:
Table 8: Performance on Language and Document Tasks (Left) Evaluation on English-Spanish and Antonym Generation tasks. (Right) MTV performance across different shot settings on document tasks.
(a) Performance on Language-Only Tasks
English-Spanish Antonym Generation
10 shot 65.2 56.0
400 shot 68.5 57.6
MTV 4-100 76.7 61.7
(b) MTV Document Task Performance
0-shot 4-shot 8-shot MTV
ChartQA 19.1 25.0 26.4 34.9
TextVQA 42.4 45.4 47.1 51.0
MTV on additional document datasets. Multimodal documents are a form of data with complex compositions of visual and textual modalities, with interleaved language, photograph, and chart information. As such, we provide some preliminary results on the effectiveness of MTV on these types of datasets in Table 8(b). These encouraging results prompt future research into the domain of leveraging task vectors for learning challenging document tasks.